Palencia Paints Corporation has a target capital structure of 45% debt and 55% common equity, with no preferred stock. Its before-tax cost of debt is 13%, and its marginal tax rate is 40%. The current stock price is P0 = $20.00. The last dividend was D0 = $2.00, and it is expected to grow at a 8% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. Do not round your intermediate calculations.

  1. rs =
  2. WACC =
Answer #1

Cost of common equity=(D1/Current price)+Growth rate

=[(2*1.08)/20]+0.08

=18.8%

After tax cost of debt=13*(1-tax rate)

=13*(1-0.4)=7.8%

WACC=Respective costs*Respective weight

=(18.8*0.55)+(7.8*0.45)

=13.85%